OpenStudy (anonymous):

This is quick, can you just check my answer please! What is y = x2 12x + 20 written in vertex form? (Points : 4) y = (x -12)^2 + 36 y = (x - 12)^2 + 26 y = (x- 6)^2 - 16 y = (x- 6)^2 + 20 I think it is C

OpenStudy (anonymous):

Oops! the question is y=x^2 -12+20

OpenStudy (dtan5457):

x2 12x + 20 +/- 12x??

OpenStudy (dtan5457):

Oh nvm

OpenStudy (dtan5457):

y=x^2-12+20 Complete the square y-20=x^2-12 (b/2)^2 for the value of c, which is 36, add to both sides y+16=x^2-12+36 y+16=(x-6)^2 subtract 16 y=(x-6)^2-16

OpenStudy (anonymous):

Vertex form=y=a(x-h)^2+k Find the vertex by x=(-b/2a) x=6 Plug it in. y=6^2-12(6)+20 y=36-72+20 y=-16 Vertex=(6,-16) y=a(x+6)^2-16 So yeah, C is right.

OpenStudy (dtan5457):

So yeah, you are right

OpenStudy (dtan5457):

epic timing @ashley_king

OpenStudy (anonymous):

ha i was just thinking that @dtan5457

OpenStudy (anonymous):

Haha thanks guys!! <3 :* Xx

OpenStudy (dtan5457):

We both have different ways of solving it though, good to look at both :)

OpenStudy (anonymous):

I don't know who to give the medal toooooo!! D: